OpenStudy (anonymous):

Sixty is what percent of 50? A. 30% B. 90% C. 110% D. 120% I got A

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Explain

OpenStudy (anonymous):

c

OpenStudy (anonymous):

use is over of so youd have \[is/of=x/100\]

OpenStudy (anonymous):

your is-60 of-50 so it would be 60/50= x/100

OpenStudy (anonymous):

ok thanks

OpenStudy (acxbox22):